Aggregator

booster cavity tuning and in SELAP

1 day 13 hours ago
Log Number: 4467610Logbook(s): Body: 

cavities were pretty far off from the machine frequency
cavity 1 - ~24 kHz
cavity 2 - ~15 kHz

tuned the cavities closer to machine frequency. increased the gradients in SEL to the set points from the last run
cavity 1 - 2.5 MV/m
cavity 2 - 10.09 MV/m

made sure drive level (GASK) and the forward power levels (CRFP) are similar to the ones from the last run

poff optimization was done after tuning the cavities. used the one button to recover the cavities into SELAP

bachiman

Operator SIGN IN

1 day 13 hours ago
Log Number: 4467609Logbook(s): Body:  ACCELERATOR OPERATOR SIGN IN FORM OPERATOR:    Smith
DATE:             01/28/2026
SHIFT:               Owl

The operator has read the following:
  • ELOG
  • PD Shift Plan
  • Radiation Surveys for PSS Segments in Controlled Access
  • PSS Log
msmith

FML SAM Update 1/29

1 day 15 hours ago
Log Number: 4467608Logbook(s): Body: 

Jan 28th
- Tunnel Crack Repairs
- Hall C Scaffolding
- NL Rack Fan Repair

Jan 29th
- Tunnel Crack Repairs
- Drip Pan Installation

nandrews

MOLLER Install Update 01/28/2026

1 day 19 hours ago
Log Number: 4467604Logbook(s): Tags: Entry Makers: jonesgaskelldwinesdougberdnikciprianBody: 

Continued installation work of LCW MOLLER system. Hall D with assistance, of Hall A techs. Craned the seven sided ring assembly onto the pion donut assembly frame. Hall A techs continued work of detaching target platform from blue utility platform.

lhurt

COMMENT ON RF HCO & Gradient Ramp 2026

1 day 20 hours ago
Log Number: 4467601Logbook(s): Body: 

Update to the No-Touch List: Cryo was experiencing issues with the heaters in 0L02, 1L08, 1L09, 1L12, and 1L21. They said operation of 0L02 is fine since the heat load is so low, but do not turn RF On in 1L08, 1L09, 1L12, and 1L21.

Additionally, when attempted to turn on HV in a zone, if it is not successful after two attempts, file an Ops-PR and move on to the next.

deir

COMMENT ON Powering Up and Performing HCO on 10 AMP & 20 Amp Trim System after SAM 2025

1 day 20 hours ago
Log Number: 4467597Logbook(s): Body: 

HCO was done on the Trim Racks in W1 service building, and all of the East Arc and Tagger service building; all magnets where set to +/- low current for 30 min and then cycled. All magnets and Trim Rack are ready for operation.

W202B17 and E301B11 has been added to our list of Trim Racks with W301B13, that we will need to investigate some issues with.

All 20 Amp Power Supplies in the East and West Arc as well as the Tagger service building, has been Powered On and place in Standby / Remote mode.

Banks, Diggs

diggs

iocsb4bcmtc restored

1 day 21 hours ago
Log Number: 4467596Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2025-09-23 17:25:52 The output from the myRestore execution was as follows: iocsb4bcmtc <NOTE> iocsb4bcmtc: Provided macro *BCM* is never used <NOTE> iocsb4bcmtc: Provided macro *BCM_CNXX* is never used Execute: /cs/op/iocs/iocsb4bcmtc/iocsb4bcmtc_myrestore.sh (success) Old : ibc2c24:bcmtcsp1_w 0 New : ibc2c24:bcmtcsp1_w 110 Old : ibc2c24:bcmtccfg_damp 1 New : ibc2c24:bcmtccfg_damp 1 Old : ibc2c24:bcmtcpidP_w 0 New : ibc2c24:bcmtcpidP_w
alarms

COMMENT ON SAM 2025 Box Supply Work

1 day 21 hours ago
Log Number: 4467594Logbook(s): Body: 

Controller chassis seems to be operating as expected. Inspected interface chassis again to find that one of the resistors apart of the voltage monitor circuit to be unsoldered. Fixed that and letting it run on control power to see if the issue is resolved.

maxwellr

iocsb4bcmtc restoration failed

1 day 21 hours ago
Log Number: 4467590Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2025-09-23 17:25:52 The output from the myRestore execution was as follows: iocsb4bcmtc <NOTE> iocsb4bcmtc: Provided macro *BCM* is never used Execute: /cs/op/iocs/iocsb4bcmtc/iocsb4bcmtc_myrestore.sh (success) Old : ibc2c24:bcmtcsp1_w 0 New : ibc2c24:bcmtcsp1_w 110 Old : ibc2c24:bcmtccfg_damp 1 New : ibc2c24:bcmtccfg_damp 1 Old : ibc2c24:bcmtcpidP_w 0 New : ibc2c24:bcmtcpidP_w 0 Old : ibc2c24:bcmtcpidI_w 0 New : ibc
alarms

COMMENT ON Removal and Installation of Girders 2L03A & 2L04A, 2L25A and 2L26A

1 day 21 hours ago
Log Number: 4467589Logbook(s): Body: 

Girder's 2L03A and 2L04A Beamline are transferred to South Linac, Removed and Reset MQB2L03 from Old Girder Plate to New Girder Plate, Girder 2L03A Beamline will be set into Girder Plate tomorrow and then setting into its resting location; we will also be setting MQB2L04 and Girder 2L04A Beamline into New Girder Plate and into its resting location.

marble

iocsb4bcmtc restoration failed

1 day 21 hours ago
Log Number: 4467588Logbook(s): Tags: Body: restorelib 3.0.1 Target restore date was 2025-09-23 17:25:52 The output from the myRestore execution was as follows: iocsb4bcmtc <NOTE> iocsb4bcmtc: Provided macro *BCM* is never used Execute: /cs/op/iocs/iocsb4bcmtc/iocsb4bcmtc_myrestore.sh (success) Old : ibc2c24:bcmtcsp1_w 0 New : ibc2c24:bcmtcsp1_w 110 Old : ibc2c24:bcmtccfg_damp 1 New : ibc2c24:bcmtccfg_damp 1 Old : ibc2c24:bcmtcpidP_w 0 New : ibc2c24:bcmtcpidP_w 0 Old : ibc2c24:bcmtcpidI_w 0 New : ibc
alarms

HYST AREA 7-0 Cycles Magnets

1 day 21 hours ago
Log Number: 4467587Logbook(s): Body: 
HYST AREA Version 7-0 has FAILED to cycle:

MQA3S08 -- E101B12 -- ea1mag -- Failed to set after 2 cycle attempts


HYST AREA Version 7-0 has successfully cycled the following:

MBC3S09V
MQA5S09
MBC5S09H
MBC5S09V
MBDBS04H
MBDBS04V
MBC7S09V
MQB1S10
MBC7S08H
MBC7S08V
MQA9S05
MBC9S05H
MBC9S05V
MQA3S09
MBC3S09H
MBT1S09V
MQA3S08
MBC3S08H
MBC3S08V
MQA5S08
MBC5S08H
MBC5S08V
MBC5S07V
MQA7S07
MBDBS03V
MBC7S07V
MQA9S04
MBC9S04H
diggs

HYST AREA 7-0 Cycles Magnets

1 day 21 hours ago
Log Number: 4467586Logbook(s): Body: 

HYST AREA Version 7-0 has successfully cycled the following:

MBT1S08H
MBT1S08V
MQA3S07
MBC3S07H
MBC3S07V
MQA5S07
MBC5S07H
MBDBS03H
MBDBS02H
MBDBS02V
MBC5S05V
MBC7S05H
MBC7S05V
MQA9S03
MBC9S03H
MBC9S03V
MQA3S06
MBT1S07V
MQA3S05
MBC3S05H
MBC3S05V
MBC5S05H

Total Time: 2.550 mins

diggs

Shift Information

Shift Swing
CEBAF Program
MCC Crew Chief Humphry
MCC Operators Janak
Program Deputy May
Announcements

Access Status

Recent PSS History

recent pss history graph